OpenStudy (anonymous):

Charlie's Computer Company charges $0.65 per pound to ship computers. Part A: Write an equation to determine the total cost, c, to ship p pounds of computers. Use your equation to determine the cost of shipping 2 pounds of computers. (6 points) Part B: If the company reduces the cost to ship computers by 0.05 per pound, write an equation to determine the total cost, c, to ship p pounds of computers with the reduced cost. (4 points)

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

hint: if one pound costs 0.65 dollars, then p pounds cost: 0.65*p

OpenStudy (anonymous):

are we doing part a?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

okay.. 0.65p ?

OpenStudy (michele_laino):

correct! \[c = 0.65 \cdot p\]

OpenStudy (anonymous):

so what do we say for part a?

OpenStudy (michele_laino):

we can say this: "the requested equation, is: \(c = 0.65 \cdot p\)"

OpenStudy (anonymous):

ok and part b?

OpenStudy (michele_laino):

now, we have to replace p with 2, so we get: \[c = 0.65 \cdot p = 0.65 \cdot 2 = ...?\]

OpenStudy (anonymous):

i dont know that one :(

OpenStudy (michele_laino):

it is simple: c=0.65*2=1.3 dollars

OpenStudy (anonymous):

oh okay.. :)

OpenStudy (michele_laino):

for part B) if one pound is shipped for 0.05 dollars, then p pounds are shipped for: 0.05*p

OpenStudy (anonymous):

what do we say for part b?

OpenStudy (michele_laino):

we can write this statement: "the cost to ship \(p\) pounds, with the reduced unitary cost is \(c=0.05 \cdot p\)"

OpenStudy (anonymous):

so whats part A and Part B... please include all of the work :)

OpenStudy (michele_laino):

summarizing, we can write this: part A) requested formula is \(c=0.65 \cdot p \) so, specializing to p=2, we have: \(c=0.65 \cdot 2=1.30$\) part B) the new formula, using the reduced unitary cost, is: \(c=0.05 \cdot p \)

OpenStudy (anonymous):

ok thanks :) i have more questions :)

OpenStudy (michele_laino):

please wait, since someone has tagged me

OpenStudy (anonymous):

okay ;)